How much pure acid must be added to 60 ml of a 35% solution to produce a mixture that is 80% acid?
ml of pure acid/ml of mixture = concentration of acid

Answer by stanbon(75887)

How much pure acid must be added to 60 ml of a 35% solution to produce a mixture that is 80% acid?
ml of pure acid/ml of mixture = concentration of acid
---------------------------
Equation:
acid + acid = acid
0.35*60 + x = 0.80(60+x)
---
Multiply thru by 100 to get:
35*60 + 100x = 80*60 + 80x
-----------------
20x = 45*60
x = 15 ml (amt. of pure acid needed)
============================
Cheers,
Stan H.